OpenStudy (phebe):

What is the slope of the line that passes through the points (-2,5) and (1,4)? A: -3 B: -2 C: -1/3 D: 1/3

OpenStudy (phebe):

@esshotwired , @magbak

OpenStudy (esshotwired):

To find the slope given two points you would use this equation: \[\frac{ x _{2}-x _{1} }{ y _{2}-y _{1} }\] So for this problem you would set it up as: \[\frac{ 1-(-2) }{ 4-5 }\] This can be simplified to solve easier. \[\frac{ 1+2 }{ 4-5 }\] Now you solve this to find the slope of the line through those two points.

OpenStudy (phebe):

ohh ok but show me more plz

OpenStudy (esshotwired):

Ok well if you solve the last fraction I gave you, it would end up that the slope is \[-\frac{ 3 }{ 1 }\] But this can be simplified to be just -3.

OpenStudy (phebe):

ok thanx sooo much

OpenStudy (esshotwired):

You're welcome :)

OpenStudy (phebe):

can you help me with another one

OpenStudy (esshotwired):

I actually have to go now. Sorry

OpenStudy (phebe):

ohh fr